# Groups with et0l co-word problem

### v1

@@ 3 Virtually free groups
Every finitely generated virtually free group is a special co-et0l-group.
We note that the fact that finitely generated virtually free groups are co-et0l-groups follows immediately from the inclusion of the context-free languages in et0l (proved in [11, Theorems 2.4,2.7]) and the fact that virtually free groups have deterministic context-free word problem, and hence context-free co-word problem. But here we prove additionally that these groups are special co-et0l-groups.
@@
Let $G$ be finitely generated and virtually free. We shall prove in Theorem 6.1 below that the property of being a special co-et0l-group does not depend on the choice of the finite generating set of the group $G$. We choose a generating set $A=B\cup T^{\pm 1}$ of $G$ in which $B$ is the inverse closure of a free generating set of a free subgroup $F$ of finite index in $G$ and $T$ is a right transversal of $F$ in $G$ with the element in $F$ removed. Then every element of $G$ can be written uniquely as a word $wt$ with $w$ a freely reduced word in $B^{*}$ and $t\in T\cup\{\epsilon\}$, where $\epsilon$ denotes the empty word. It was shown in [13] that the word problem $\mathsf{WP}(G,A)$ is the language of a deterministic pushdown automaton with states corresponding to $T\cup\{\epsilon\}$, that stores the freely reduced word $w$ on its stack. We can amend that to a pushdown automaton $\mathcal{P}_{G}$ that accepts by empty stack, by storing $t$ (if non-empty) on the stack above $w$.
We convert $\mathcal{P}_{G}$ to a cspda by attaching a check stack onto which we write a string ${\sf p}^{k}$ at the start, for some arbitrary integer $k>0$. After the start, the cspda then operates as $\mathcal{P}_{G}$, additionally reporting failure if at any stage the pushdown stack contains more than $k$ cells (and so the top of the pushdown stack is above the top of the check stack). The new automaton accepts if and only if $\mathcal{P}_{G}$ does not accept.
It is straightforward to verify that this cspda is special. Note that failure can only arise as a result of inadequate initialisation of the check stack, and this can be remedied without otherwise affecting its operation by pushing more p symbols onto the check stack. ∎
@@ 4 Bounded automata groups
This section is devoted to the proof of the following result, which extends the main result of [2]; its proof is based on the proof of that result and the proof of the corresponding result in  [9] on which it is based.
@@
Every finitely generated bounded automata group is a special co-et0l-group.
@@
For $d\geq 2$, we consider the set $\Sigma^{*}$ of strings over the finite set $\Sigma$ of $d$ elements, and write $\epsilon$ for the empty string. There is a natural bijection between $\Sigma^{*}$ and the set of vertices of the $d$-regular rooted tree $\mathcal{T}_{\Sigma}$, in which the $d$ edges down from any given vertex are labelled by the elements of $\Sigma$; similarly there is a natural bijection between $\Sigma^{\omega}$ and the set of ends of $\mathcal{T}_{\Sigma}$. We consider the group of all permutations of $\Sigma^{*}$ that induce automorphisms of $\mathcal{T}_{\Sigma}$, which we shall denote here by $\mathcal{A}_{\Sigma}$ (it is denoted by $\mathsf{Aut}(\mathcal{T}_{d})$ in [2]); such a permutation $\alpha$ must preserve ‘levels’ in the tree, and hence for any $k$ it permutes the set of subsets of strings with common prefixes of length $k$; that is the subsets $\mathbf{x}\Sigma^{*}$ of $\Sigma^{*}$ with $\mathbf{x}\in\Sigma^{k}$. In addition it must have the property that, for any $\mathbf{x}\in\Sigma^{*}$, $\mathbf{y}\in\Sigma^{*}\cup\Sigma^{\omega}$, $(\mathbf{x}\mathbf{y})^{\alpha}$ has $\mathbf{x}^{\alpha}$ as a prefix, and the restriction $\alpha|_{\mathbf{x}}$ of $\alpha$ to $\mathbf{x}$ is defined by $(\mathbf{x}\mathbf{y})^{\alpha}=\mathbf{x}^{\alpha}\mathbf{y}^{\alpha|_{\mathbf{x}}}$.
@@
An automorphism $\phi\in\mathcal{A}_{\Sigma}$ is called finitary if, for some $k_{\phi}\in{\mathbb{N}}$ and all $\mathbf{x}$ with $|\mathbf{x}|\geq k_{\phi}$, the restriction $\phi|_{\mathbf{x}}$ is equal to the identity. We refer the reader to [2] for the precise definition of a directed automaton automorphism; such an automorphism $\delta$ has an associated spine, an element of $\Sigma^{\omega}$, with the property that $\delta|_{\mathbf{y}}$ is finitary for any $\mathbf{y}$ that is not a prefix of $\mathsf{spine}(\delta)$. It is proved in [2, Lemma 1] that the spine of any directed automaton automorphism is eventually periodic, that is, of the form $\mathbf{p}\mathbf{q}^{\omega}$, for fixed strings $\mathbf{p},\mathbf{q}\in\Sigma^{*}$, and that $(\mathbf{p}\mathbf{q}^{\omega})^{\delta}=\mathbf{p}^{\prime}\mathbf{q}^{\prime\omega}$, for fixed strings $\mathbf{p}^{\prime},\mathbf{q}^{\prime}\in\Sigma^{*}$ that have the same lengths as $\mathbf{p}$ and $\mathbf{q}$. Furthermore, if $\mathbf{q}=b_{1}\cdots b_{t}$ with $b_{i}\in\Sigma$, then $\delta|_{\mathbf{p}\mathbf{q}^{k}b_{1}\cdots b_{j}}=\delta|_{\mathbf{p}b_{1}\cdots b_{j}}$ for any $k\geq 0$ and $0\leq j<t$.
Our proof of Theorem 4.1 (below) is a modification of the proof of [2], which itself emulates the proof in [9]; We shall describe a construction that is basically the construction of [2], with some modification to fit the constraints that we have imposed on a cspda to make it special. We shall not repeat proofs of properties of the cspda that are already proved in [2].
@@
Proof of Theorem 4.1. Let $G$ be a subgroup of $\mathcal{D}_{\Sigma}$ with finite generating set $X$.
It is proved in [9, Proposition 16] that $\mathcal{D}_{\Sigma}$ is generated by the infinite subset of all of its finitary and directed automaton automorphisms. Hence the subgroup $G$ of $\mathcal{D}_{\Sigma}$ is a finitely generated subgroup of a group with finite generating set $Y$, each of whose elements is either finitary or a directed automaton automorphism. It will follow from Theorem 6.1 below that $G$ is a special co-et0l group provided that the subgroup $\langle Y\rangle$ of $\mathcal{D}_{\Sigma}$ is special co-et0l. Hence, as in [2], we may assume that $G$ itself has such a generating set $Y$.
@@
Stage 1 of cspda that we define here is the same as in [2], with $x_{1}$ as the top symbol on the check stack. In the construction of [2], $\mathbf{x}$ is put immediately onto the pushdown stack at the end of Stage 1 but, in our modification, to ensure that our cspda is special, we defer copying the contents of the check stack onto the pushdown stack until after we have entered an entry state of the cspda and read the first letter in the input word. This is because Property (1) of a special cspda requires that the pushdown stack be empty when the first input symbol is read.
@@
Our Stage 2 process is similar to that of [2], but it does a little more than in [2]. For, in order to maintain determinism after Stage 1, we do not have a Stage 3, but do our checking within Stage 2; the checking will be explained later. At Stage 2, we have just two reading states. The first is the unique entry state and is non-accepting, and the second is an accepting state; we shall call those two states $q_{1}$ and $q_{2}$ respectively. All non-reading states are non-accepting.
@@
where in both cases $\phi$ is the finitary automorphism $\delta|_{\mathbf{z}y_{\ell+1}}$ with associated bound $k_{\phi}$ on the length of the prefix of words that it moves, and $\ell_{\phi}:=\min(\ell+1+k_{\phi},m)$.
@@
The case where $\ell\leq s$ is straightforward. The computation and storage of $y_{\ell+1}^{\delta_{\mathbf{z}}}$ and $(y_{\ell+2},\cdots,y_{\ell_{\phi}})^{\phi}$ need a bounded amount of memory, and so can be done deterministically using non-reading states and transitions of the cspda. Then we push $((y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi})^{\mathsf{R}}$ onto the pushdown stack followed by $(y_{\ell+1})^{\delta|_{\mathbf{z}}}a^{\prime}_{\ell}\cdots a^{\prime}_{1}$.
The case where $\ell\geq s$ is more complicated. As explained above, it is proved in [2, Lemma 1] that $\delta|_{\mathbf{z}}=\delta|_{\mathbf{p}\mathbf{q}^{k}b_{1}\cdots b_{j}}$ does not depend on $k$, and so $\delta|_{\mathbf{z}}$ and the finitary automorphism $\phi$ depend only on $\mathbf{p}$, $\mathbf{q}$ and $j$. So the computation and storage of the components of $\mathbf{y}^{\delta}$, namely $\mathbf{p}^{\prime}$, $\mathbf{q}^{\prime}$, $y_{\ell+1}^{\delta_{\mathbf{z}}}$ and $(y_{\ell+2},\cdots,y_{\ell_{\phi}})^{\phi}$, is again straightforward, and needs only a bounded amount of memory. As above, we can then push $((y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi})^{\mathsf{R}}$ onto the pushdown stack followed by $(y_{\ell+1})^{\delta|_{\mathbf{z}}}b^{\prime}_{j}\cdots b^{\prime}_{1}$ But then we need to push $k$ copies of $\mathbf{q}^{\prime\mathsf{R}}$ followed by $\mathbf{p}^{\prime\mathsf{R}}$ without knowing the value of $k$ in advance. And $k$ could take any non-negative integer value. This computation cannot be achieved using bounded memory.
@@
In the process described in [2], the cspda decides after processing each input symbol either to read another input symbol or to proceed to Stage 3, and look for no further input. If the machine moves to Stage 3 prematurely, then the end of the input word is never reached, and so that word is not accepted. If the machine looks for input after reading the end of the word, then it must also fail. The machine can only accept the word if it makes that transition at the right moment and finds that the contents of the two stacks are distinct. Clearly this part of the process of [2] is non-deterministic.
@@
It is proved in [2] that the machine described there is a cspda that accepts an input word precisely when it is in $\mathsf{coWP}(G,A)$. It should be clear that our modified cspda accepts the same language as does the cspda of [2], and that our modifications ensure that properties (0), (1) and (2) hold for our cspda.
It remains to consider property (3). Since our machine never fails, we only need to check (3) (a), that for any words $u\in A^{*},v\in\mathsf{WP}(G,A)$ the configurations $C^{u}$ and $C^{uv}$ are equal. These configurations are determined by the state of the machine and the contents of the pushdown stack. At any stage the pushdown stack must contain the image of the string on the check stack under the action of the element represented by the input string that has been read so far. Since $u,uv$ represent the same group element, the pushdown stack contents in the two configurations must match, as required. And the state within each configuration must be either $q_{1}$ or $q_{2}$, depending on whether or not that single element is equal to the identity. $\Box$
@@ 5 Higman–Thompson groups
The Higman–Thompson groups $G_{n,r}$ groups were described by Higman in [8] as generalizations of one of Thompson’s finitely presented infinite simple groups, namely the group $V$ described in [15], which is isomorphic to $G_{2,1}$, and which first appeared in a handwritten manuscript of Richard J. Thompson circa 1965. They were proved to be co-indexed in [9] and the stronger result that they are co-context-free is proved in [12]. In this section we extend the proof of [9] to derive the following result.
@@
The Higman–Thompson groups $G_{n,r}$ are special co-et0l-groups.
@@
There are several equivalent definitions of the groups $G_{n,r}$, and the one we use here is taken from [10, Section 1.10.11], and is also used in [9, Section 3]. Our proof that these groups are special co-et0l is based on the proof in [9, Theorem 1] that they are stack groups, and is similar to our proof of Theorem 4.1.
Fix integers $r\geq 1$ and $n\geq 2$, and let $Q:=\{q_{1},q_{2},\ldots,q_{r}\}$ and $\Sigma:=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{n}\}$ be disjoint sets of cardinality $r$ and $n$ respectively. Define $\Omega$ to be the set of all infinite strings of the form $\omega=q_{i}\sigma_{i_{1}}\sigma_{i_{2}}\sigma_{i_{3}}\cdots$, where $q_{i}\in Q$ and $\sigma_{i_{j}}\in\Sigma$ for $j\geq 1$. A non-empty prefix of $\omega$ is a finite string (or word) $q_{i}\sigma_{i_{1}}\cdots\sigma_{i_{k}}$ for some $k\geq 0$. Roughly speaking $G_{n,r}$ is the group which acts on $\Omega$ by replacing non-empty prefixes; we give more detail below.
@@
A bijection $\varphi:B\to C$ between two finite complete antichains induces a permutation of $\Omega$, also called $\varphi$, defined by $\omega^{\varphi}:=b^{\varphi}w^{\prime}$, where $\omega=b\omega^{\prime}$ and $b$ is the unique prefix of $\omega$ that lies in $B$. The group $G_{n,r}$ is defined to be the group of all such permutations of $\Omega$. We also define $\mathbf{x}^{\varphi}:=b^{\varphi}\omega^{\prime}$ for any prefix $\mathbf{x}=b\omega^{\prime}$ of an element of $\Omega$ that has $b\in B$ as a prefix.
Proof of Theorem 5.1. This proof is very similar to that of Theorem 4.1 and the computation proceeds by making prefix replacements to a word stored on the pushdown stack. The main difference is that prefixes may be replaced by words of a different length. As in the earlier proof, there are just two reading states $q_{1}$ and $q_{2}$, which are respectively non-accepting and accepting, where the pushdown stack is empty in state $q_{1}$ and non-empty in state $q_{2}$. All non-reading states are non-accepting.
@@
In its first stage, the cspda with language $\mathsf{coWP}(G,A)$ puts an arbitrary word $\mathbf{x}$ in $Q\Sigma^{*}$ on the check stack, with its first symbol $q_{i}$ at the top, and then puts an arbitrary number of padding symbols on top of that. The read write head then moves to the bottom of the stacks, and the cspda enters its deterministic second stage in the non-accepting reading state $q_{1}$ with the pushdown stack empty. After reading the first letter (if any) of the input word $\alpha$, the word $\mathbf{x}$ is copied onto the pushdown stack. Thereafter, the cspda works by replacing $\mathbf{x}$ by $\mathbf{x}^{\beta}$, where $\beta$ is the prefix of $\alpha$ that has been read so far.
The cspda attempts to establish that the input word $\alpha$ does not fix some string in $\Omega$ with prefix $\mathbf{x}$, which would prove that it lies in $\mathsf{coWP}(G,A)$. To do that, it attempts to compute the image $\mathbf{x}^{\alpha}$ of $\mathbf{x}$ under $\alpha$, and then to check whether $\mathbf{x}^{\alpha}$ is unequal to $\mathbf{x}$, which would prove that $\alpha\in\mathsf{coWP}(G,A)$. If $\alpha$ does lie in $\mathsf{coWP}(G,A)$, then there exists a string in $\Omega$ that it does not fix and, provided that a sufficiently long prefix $\mathbf{x}$ of this string together with enough padding symbols are put on the check stack in the first stage, the cspda will succeed in showing that $w\not\in\mathsf{coWP}(G,A)$.
@@
As in the proof of Theorem 4.1, before entering a reading state, the cspda checks whether the words $\mathbf{x}^{\beta}$ and $\mathbf{x}$ on the stacks are equal. If they have different lengths or, equivalently, if the letter that the read write head sees on the check stack does not lie in the set $Q$, then the words cannot be equal, and the cspda moves immediately into the accepting state $q_{2}$, since there is an element of $\Omega$ that is not fixed. Otherwise, as in the proof of Theorem 4.1, the cspda checks the words for equality by emptying the pushdown stack of letters that are the same as the corresponding letters on the check stack. It then goes into state $q_{1}$ if the pushdown stack is empty, and into $q_{2}$ if it is non-empty. After reading the next input letter if any, if the pushdown stack is empty, or if the symbol currently at its top does not lie in $Q$, then the cspda reconstitutes the word $\mathbf{x}^{\beta}$ on the pushdown stack by copying letters from the check stack.
@@
6 Closure properties of special co-et0l-groups
Our proofs of the closure properties in this section closely follow the proofs of corresponding properties for stack groups in [9, Sections 7 and 8].

### v2

@@ 3 Virtually free groups
Theorem 3.1, which we prove in this section, is the special case of Theorem 6.2 in which the group $H$ is trivial. But we prefer to prove Theorem 3.1 first, firstly because it is more straightforward, and secondly because the technique used in the proof is applied again in the proof of Theorem 6.2. We observe in addition that, since the Thompson group $V$ is known to contain all virtually free groups as subgroups (see, for example, [4, Lemma 1.1, Theorem 1.4, Corollary 1.3]), this result could also be deduced from Theorems 5.1 and 6.1.
We note that in the proof of Theorem 3.1, and also in the proofs of Theorems 4.1 and 5.1, we refer forwards to and use Theorem 6.1. The proof of that result is elementary, and does not depend on earlier results in this article, and we have chosen to defer its proof to Section 6, which is devoted to closure properties of the class of special co-et0l groups.
@@
Every finitely generated virtually free group is a special co-et0l group.
We note that the fact that finitely generated virtually free groups are co-et0l groups follows immediately from the inclusion of the context-free languages in et0l (proved in [13, Theorems 2.4,2.7]) and the fact that virtually free groups have deterministic context-free word problem, and hence context-free co-word problem. But here we prove additionally that these groups are special co-et0l groups, by constructing a special cspda.
@@
Let $G$ be finitely generated and virtually free. We shall prove in Theorem 6.1 below that the property of being a special co-et0l group does not depend on the choice of the finite generating set of the group $G$. We choose a generating set $A=B\cup T^{\pm 1}$ of $G$ in which $B$ is the inverse closure of a free generating set of a free subgroup $F$ of finite index in $G$ and $T$ is a right transversal of $F$ in $G$ with the element in $F$ removed. Then every element of $G$ can be written uniquely as a word $wt$ with $w$ a freely reduced word in $B^{*}$ and $t\in T\cup\{\epsilon\}$, where $\epsilon$ denotes the empty word. It was shown in [15] that the word problem $\mathsf{WP}(G,A)$ is the language of a deterministic pushdown automaton with states corresponding to $T\cup\{\epsilon\}$, that stores the freely reduced word $w$ on its stack. We can amend that to a pushdown automaton $\mathcal{P}_{G}$ that accepts when its stack is empty, by storing $t$ (if non-empty) on the stack above $w$.
We convert $\mathcal{P}_{G}$ to a cspda by attaching a check stack onto which we write a string $\mathfrak{b}{\sf p}^{k}\mathfrak{t}$ at the start, for some arbitrary integer $k>0$. It has just two states, the entry state, which is entered if and only if $\mathfrak{b}$ is seen on the two stacks, and an accepting state. After the initialisation and writing $\mathfrak{b}$ on the pushdown stack, the cspda operates as $\mathcal{P}_{G}$, additionally reporting failure if (P3) is violated by an attempt to push a symbol onto the pushdown stack above $\mathfrak{t}$. The new automaton accepts any word that $\mathcal{P}_{G}$ rejects, provided that $k$ has been chosen to be large enough, and rejects if $\mathcal{P}_{G}$ accepts (or fails if $k$ is too small).
It is straightforward to verify that this cspda is special with language $\mathsf{coWP}(G,A)$. Note that failure can only arise as a result of inadequate initialisation of the check stack, and this can be remedied without otherwise affecting its operation by pushing more p symbols onto the check stack. ∎
@@ 4 Bounded automata groups
This section is devoted to the proof of the following result, which strengthens the main result of [2]; its proof is based on the proof of that result and the proof of the corresponding result in [11] on which it is based.
@@
Every finitely generated bounded automata group is a special co-et0l group.
@@
For $d\geq 2$, we consider the set $\Sigma^{*}$ of strings over the finite set $\Sigma$ of $d$ elements, and write $\epsilon$ for the empty string. There is a natural bijection between $\Sigma^{*}$ and the set of vertices of the $d$-regular rooted tree $\mathcal{T}_{\Sigma}$, in which the $d$ edges down from any given vertex are labelled by the elements of $\Sigma$; similarly there is a natural bijection between $\Sigma^{\omega}$ and the set of ends of $\mathcal{T}_{\Sigma}$. We consider the group of all permutations of $\Sigma^{*}$ that induce automorphisms of $\mathcal{T}_{\Sigma}$, which we shall denote here by $\mathcal{A}_{\Sigma}$ (it is denoted by $\mathsf{Aut}(\mathcal{T}_{d})$ in [2]); such a permutation $\alpha$ must preserve ‘levels’ in the tree, and hence for any $k$ it permutes the set of subsets of strings with common prefixes of length $k$; that is the subsets $\mathbf{x}\Sigma^{*}$ of $\Sigma^{*}$ with $\mathbf{x}\in\Sigma^{k}$. In addition it must have the property that, for any $\mathbf{x}\in\Sigma^{*}$ and, $\mathbf{y}\in\Sigma^{*}\cup\Sigma^{\omega}$, the string $(\mathbf{x}\mathbf{y})^{\alpha}$ has $\mathbf{x}^{\alpha}$ as a prefix, and the restriction $\alpha|_{\mathbf{x}}$ of $\alpha$ to $\mathbf{x}$ is defined by $(\mathbf{x}\mathbf{y})^{\alpha}=\mathbf{x}^{\alpha}\mathbf{y}^{\alpha|_{\mathbf{x}}}$.
@@
An automorphism $\phi\in\mathcal{A}_{\Sigma}$ is called finitary if, for some $k_{\phi}\in{\mathbb{N}}$ and all $\mathbf{x}$ with $|\mathbf{x}|\geq k_{\phi}$, the restriction $\phi|_{\mathbf{x}}$ is equal to the identity.
We refer the reader to [2] for the precise definition of a directed automaton automorphism; such an automorphism $\delta$ has an associated spine, an element of $\Sigma^{\omega}$, with the property that $\delta|_{\mathbf{y}}$ is finitary for any $\mathbf{y}$ that is not a prefix of $\mathsf{spine}(\delta)$. It is proved in [2, Lemma 1] that the spine of any directed automaton automorphism is eventually periodic, that is, of the form $\mathbf{p}\mathbf{q}^{\omega}$, for fixed strings $\mathbf{p},\mathbf{q}\in\Sigma^{*}$, and that $(\mathbf{p}\mathbf{q}^{\omega})^{\delta}=\mathbf{p}^{\prime}\mathbf{q}^{\prime\omega}$, for fixed strings $\mathbf{p}^{\prime},\mathbf{q}^{\prime}\in\Sigma^{*}$ that have the same lengths as $\mathbf{p}$ and $\mathbf{q}$. Furthermore, $\mathbf{p}$ and $\mathbf{q}$ can be chosen such that, if $\mathbf{q}=b_{1}\cdots b_{t}$ with $b_{i}\in\Sigma$, then $\delta|_{\mathbf{p}\mathbf{q}^{k}b_{1}\cdots b_{j}}=\delta|_{\mathbf{p}b_{1}\cdots b_{j}}$ for any $k\geq 0$ and $0\leq j<t$.
Our proof of Theorem 4.1 (below) is a modification of the proof of [2], which itself emulates the proof in [11]. We shall describe a construction that is basically the construction of [2], with some modification to fit the constraints that we have imposed on a cspda to make it special. We shall not repeat proofs of properties of the cspda that are already proved in [2].
@@
Proof of Theorem 4.1. Let $G$ be a subgroup of $\mathcal{D}_{\Sigma}$ with finite generating set $X$. It is proved in [11, Proposition 16] that $\mathcal{D}_{\Sigma}$ is generated by the infinite subset of all of its finitary and directed automaton automorphisms. Hence the subgroup $G$ of $\mathcal{D}_{\Sigma}$ is a finitely generated subgroup of a group with finite generating set $Y$, each of whose elements is either finitary or a directed automaton automorphism. It will follow from Theorem 6.1 below that $G$ is a special co-et0l group provided that the subgroup $\langle Y\rangle$ of $\mathcal{D}_{\Sigma}$ is special co-et0l. Hence, as in [2], we may assume that $G$ itself has such a generating set $Y$.
@@
Stage 1 of the cspda that we define here is essentially the same as in [2], with $x_{1}$ as the top symbol on the check stack, but to satisfy Property (P3) we also write $\mathfrak{t}$ on the check stack above $x_{1}$. In the construction of [2], $\mathbf{x}$ is put immediately onto the pushdown stack at the end of Stage 1 but, in our modification, to ensure that (P3) is satisfied, we defer copying the contents of the check stack onto the pushdown stack until after we have entered an entry configuration of the cspda and then have read the first letter in the input word.
@@
Our Stage 2 process is similar to that of [2], but it does a little more than in [2]. For, in order to maintain determinism after Stage 1, we do not have a Stage 3, but do our checking within Stage 2; the checking will be explained later. In our Stage 2, we have just two reading states. The first is the unique entry state and is non-accepting, and the second is an accepting state; we shall call those two states $q_{1}$ and $q_{2}$, respectively. As required by (P5), all non-reading states are non-accepting.
@@
where in both cases $\phi$ is the finitary automorphism $\delta|_{\mathbf{z}y_{\ell+1}}$ with associated bound $k_{\phi}$ on the length of the prefix of words that it moves, and $\ell_{\phi}\coloneqq\min(\ell+1+k_{\phi},m)$.
@@
The case where $\ell\leq s$ is straightforward. The computation and storage of $y_{\ell+1}^{\delta_{\mathbf{z}}}$ and $(y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi}$ need a bounded amount of memory, and so can be done deterministically using non-reading states and transitions of the cspda. Then we push $((y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi})^{\mathsf{R}}$ onto the pushdown stack followed by $(y_{\ell+1})^{\delta|_{\mathbf{z}}}a^{\prime}_{\ell}\cdots a^{\prime}_{1}$.
The case where $\ell\geq s$ is more complicated. As explained above, it is proved in [2, Lemma 1] that $\delta|_{\mathbf{z}}=\delta|_{\mathbf{p}\mathbf{q}^{k}b_{1}\cdots b_{j}}$ does not depend on $k$, and so $\delta|_{\mathbf{z}}$ and the finitary automorphism $\phi$ depend only on $\mathbf{p}$, $\mathbf{q}$ and $j$. So the computation and storage of the components of $\mathbf{y}^{\delta}$, namely $\mathbf{p}^{\prime}$, $\mathbf{q}^{\prime}$, $y_{\ell+1}^{\delta_{\mathbf{z}}}$ and $(y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi}$, is again straightforward, and needs only a bounded amount of memory. As above, we can then push $((y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi})^{\mathsf{R}}$ onto the pushdown stack followed by $(y_{\ell+1})^{\delta|_{\mathbf{z}}}b^{\prime}_{j}\cdots b^{\prime}_{1}$ But then we need to push $k$ copies of $\mathbf{q}^{\prime\mathsf{R}}$ followed by $\mathbf{p}^{\prime\mathsf{R}}$ without knowing the value of $k$ in advance, and $k$ could take any non-negative integer value. This computation cannot be achieved using bounded memory.
@@
In the process described in [2], the cspda decides after processing each input symbol either to read another input symbol, or to proceed to Stage 3 and look for no further input. If the machine moves to Stage 3 prematurely, then the end of the input word is never reached, and so that word is not accepted. If the machine looks for input after reading the end of the word, then it must also fail. The machine can only accept the word if it makes that transition at the right moment and finds that the contents of the two stacks are distinct. Clearly this part of the process of [2] is non-deterministic.
@@
It is proved in [2] that the machine described there is a cspda that accepts an input word precisely when it is in $\mathsf{coWP}(G,A)$. It should be clear that our modified cspda accepts the same language as does the cspda of [2], and that our modifications ensure that properties (P1)–(P5) and (P7) hold for our cspda.
It remains to consider Property (P6). Since our machine never fails, we only need to check (P6) (a), that for any words $u\in A^{*},v\in\mathsf{WP}(G,A)$ the configurations $C^{u}$ and $C^{uv}$ are equal. These configurations are determined by the state of the machine and the contents of the pushdown stack. At any stage the pushdown stack must contain the image of the string on the check stack under the action of the element represented by the input string that has been read so far. Since $u,uv$ represent the same group element, the pushdown stack contents in the two configurations must match, as required. And the state within each configuration must be either $q_{1}$ or $q_{2}$, depending on whether or not that single element is equal to the identity. $\Box$
@@ 5 Higman–Thompson groups
The Higman–Thompson groups $G_{n,r}$ were described by Higman in [10] as generalisations of one of Thompson’s finitely presented infinite simple groups, namely the group $V$ described in [17], which is isomorphic to $G_{2,1}$, and which first appeared in a handwritten manuscript of Richard J. Thompson circa 1965. They were proved to be co-indexed in [11] and the stronger result that they are co-context-free is proved in [14]. In this section we extend the proof of [11] to derive the following result.
@@
The Higman–Thompson groups $G_{n,r}$ are special co-et0l groups.
@@
There are several equivalent definitions of the groups $G_{n,r}$, and the one we use here is taken from [12, Section 1.10.11], and is also used in [11, Section 3]. Our proof that these groups are special co-et0l is based on the proof in [11, Theorem 1] that they are stack groups, and is similar to our proof of Theorem 4.1.
Fix integers $r\geq 1$ and $n\geq 2$, and let $Q\coloneqq\{q_{1},q_{2},\ldots,q_{r}\}$ and $\Sigma\coloneqq\{\sigma_{1},\sigma_{2},\ldots,\sigma_{n}\}$ be disjoint sets of cardinality $r$ and $n$ respectively. Define $\Omega$ to be the set of all infinite strings of the form $\omega=q_{i}\sigma_{i_{1}}\sigma_{i_{2}}\sigma_{i_{3}}\cdots$, where $q_{i}\in Q$ and $\sigma_{i_{j}}\in\Sigma$ for $j\geq 1$. A non-empty prefix of $\omega$ is a finite string (or word) $q_{i}\sigma_{i_{1}}\cdots\sigma_{i_{k}}$ for some $k\geq 0$. Roughly speaking $G_{n,r}$ is the group which acts on $\Omega$ by replacing non-empty prefixes; we give more detail below.
@@
A bijection $\varphi\colon B\to C$ between two finite complete antichains induces a permutation of $\Omega$, also called $\varphi$, defined by $\omega^{\varphi}\coloneqq b^{\varphi}w^{\prime}$, where $\omega=b\omega^{\prime}$ and $b$ is the unique prefix of $\omega$ that lies in $B$. The group $G_{n,r}$ is defined to be the group of all such permutations of $\Omega$. We also define $\mathbf{x}^{\varphi}\coloneqq b^{\varphi}\omega^{\prime}$ for any prefix $\mathbf{x}=b\omega^{\prime}$ of an element of $\Omega$ that has $b\in B$ as a prefix.
Proof of Theorem 5.1. This proof is very similar to that of Theorem 4.1 and the computation proceeds by making prefix replacements to a word stored on the pushdown stack. The main difference is that prefixes may be replaced by words of a different length. As in the earlier proof, there are just two reading states $q_{1}$ and $q_{2}$, which are respectively non-accepting and accepting, where the state is $q_{1}$ when the pushdown stack contains only $\mathfrak{b}$, and $q_{2}$ otherwise. All non-reading states are non-accepting (as required by (P5)).
@@
In its first stage, the cspda with language $\mathsf{coWP}(G,A)$ puts an arbitrary word $\mathbf{x}$ in $Q\Sigma^{*}$ on the check stack, with its first symbol $q_{i}$ at the top, and then puts an arbitrary number of padding symbols on top of that followed by $\mathfrak{t}$. The read write head then moves to the bottom of the stacks, and the cspda enters its deterministic second stage, puts $\mathfrak{b}$ on its pushdown stack, and goes into its non-accepting reading state $q_{1}$. After the first letter (if any) of the input word $\alpha$ has been read, the word $\mathbf{x}$ is copied onto the pushdown stack. Thereafter, the cspda works by replacing $\mathbf{x}$ by $\mathbf{x}^{\beta}$, where $\beta$ is the prefix of $\alpha$ that has been read so far.
The cspda attempts to establish that the input word $\alpha$ does not fix some string in $\Omega$ with prefix $\mathbf{x}$, which would prove that it lies in $\mathsf{coWP}(G,A)$. To do that, it attempts to compute the image $\mathbf{x}^{\alpha}$ of $\mathbf{x}$ under $\alpha$, and then to check whether $\mathbf{x}^{\alpha}$ is distinct from $\mathbf{x}$, which would prove that $\alpha\in\mathsf{coWP}(G,A)$. If $\alpha$ does lie in $\mathsf{coWP}(G,A)$, then there exists a string in $\Omega$ that it does not fix and, provided that a sufficiently long prefix $\mathbf{x}$ of this string together with enough padding symbols are put on the check stack in the first stage, the cspda will succeed in showing that $w\notin\mathsf{coWP}(G,A)$.
@@
As in the proof of Theorem 4.1, before entering a reading state, the cspda checks whether the words $\mathbf{x}^{\beta}$ and $\mathbf{x}$ on the stacks are equal. If they have different lengths or, equivalently, if the letter that the read write head sees on the check stack does not lie in the set $Q$, then the words cannot be equal, and we cannot have $\beta=_{G}1$. To see this note that, for example, the infinite string $\mathbf{x}\sigma_{1}\sigma_{2}\sigma_{1}\sigma_{2}^{2}\sigma_{1}\sigma_{2}^{3}\cdots$ is not fixed by $\beta$. In this case the cspda moves immediately into the accepting state $q_{2}$.
Otherwise, as in the proof of Theorem 4.1, the cspda checks the words for equality by emptying the pushdown stack of letters that are the same as the corresponding letters on the check stack. It then goes into state $q_{1}$ if the pushdown stack contains only $\mathfrak{b}$, and into $q_{2}$ otherwise. After reading the next input letter if any, if the pushdown stack contains only $\mathfrak{b}$, or if the symbol currently at its top does not lie in $Q$, then the cspda reconstitutes the word $\mathbf{x}^{\beta}$ on the pushdown stack by copying letters from the check stack.
@@
6 Closure properties of special co-et0l groups
Our proofs of the closure properties in this section closely follow the proofs of corresponding properties for stack groups in [11, Sections 7 and 8].